Solve the following pair of linear equations by the substitution method
 
√2 x+√3 y = 0
√3 x-√8 y = 0

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student


Given,
√2 x + √3 y = 0
and
√3 x - √8 y = 0
are the two equations.
From 1st equation,
we get,
x = - (√3/√2)y __________________(1)
Putting the value of x in the given second equation to get,
√3(-√3/√2)y - √8y = 0
⇒ (-3/√2)y- √8 y = 0
⇒ y = 0

Now, substitute the value of y in equation (1),
we get,
x = 0
Therefore, x = 0 and y = 0 is required solution.

Thanks
surya prakash
3 years ago
Welcome to askIITians,
 
Given equations are  
√2 x + √3 y = 0    …..........(1)
√3 x - √8 y = 0     …..........(2)
from equation(2)  ,   y =  √3 x /  √8
Put the value of y in  equation (1) then
√3 x /  √8) =0√2 x + √3 ( 
√2 x +  3 x /  √8 =0
(16 x + 3x) /   √8 =0
 
19 x = 0   then x=0
 
put x = 0 in equation (1) 
 
√2 (0) + √3 y = 0  then y =0
we have solution set is x = 0 and y =0
 
Thanks